PK Berhad purchased a Truck for RM85, 000 on 1st of January 2013. It was paid by cash with a trade-in of RM15, 000 on an older lorry. The lorry was purchased on the 1st of January 2009 for RM70, 000 and expected to have a salvage value of RM5, 000 at the end of 5 years of its useful life.

Required:

i. Calculate the depreciation per annum on the lorry. (3 marks)


ii. Calculate the net book value of the lorry as at 31st December 2012. (6 marks)


iii. Record the journal entry for the purchase of the truck. (6 marks)
1
Expert's answer
2014-04-21T09:28:02-0400
i. Calculate the depreciation per annum on the lorry.
Depreciation per annum = (purchase value - salvage value)/number of years = (70,000 - 5,000)/5 = RM13,000
ii. Calculate the net book value of the lorry as at 31st December 2012. (6 marks)
net book value = purchase value - depreciation per annum*number of years = 70,000 - 13,000*4 = RM18,000
iii. Record the journal entry for the purchase of the truck. (6 marks)
If you purchase a $85,000 the truck with cash, the journal entry that you use to record this purchase debits delivery truck for $85,000 and credits cash for $85,000.